OpenStudy (cutiecomittee123):

What is the domain of y=ln(x+3)+2? also how do i find the domain of any logarithmic function in general.

OpenStudy (solomonzelman):

you can only take a log of something that is greater than 0, and therefore `x+3` has to be greater than 0. x+3>0 now, solve for x./

OpenStudy (cutiecomittee123):

and that will give me the domain?

OpenStudy (solomonzelman):

yes, if you solve it correctly

OpenStudy (cutiecomittee123):

I get x>-3

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

that means that x can be any number greater than -3.

OpenStudy (cutiecomittee123):

Okay thanks. is that it? so its just all numbers bigger than -3?

OpenStudy (solomonzelman):

in this case, yes/

OpenStudy (solomonzelman):

you want a little general review too, or not?

OpenStudy (cutiecomittee123):

Sure:)

OpenStudy (cutiecomittee123):

Also can you help me to find the range???

OpenStudy (solomonzelman):

The domain of a logarithmic function in general: \(\large\color{black}{ \displaystyle {\rm y} =\log_{\rm a}\left(~ {\rm f}(x)~\right)}\) above \(\color{blue}{\LARGE\Uparrow}\) is a logarithmic function. (it is a y, function of x. It has some positive base a, and any involving x function f(x) the log of which you are taking.) in this case the domain is the x-solutions that make f(x) be greater than 0. (that is pretty much all)

OpenStudy (solomonzelman):

however the range would be \((-\infty,~+\infty)\) (this is also a good place to introduce the concept of a "limit" )

OpenStudy (cutiecomittee123):

Okay well how do you find the range using the equation that I had?

OpenStudy (solomonzelman):

have you previously been finding the range of such (logarithmic) functions?

OpenStudy (solomonzelman):

I will start from a pretty simple example

OpenStudy (cutiecomittee123):

I think I may have done things like this last year when I was in algebra 2 but now i am in precalc and I dont remember how to find the range or domain of a logarithmic function.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle {\rm y}=\log(x) }\) (when the base of a log is not specified, the rule is that this base is 10) Obviously, as x approaches \(\infty\) (saying - gets infinitely bigger and bigger), the output [in this case] is also getting larger and larger. So, we can tell that as \(x\rightarrow \infty\), then \(y\rightarrow \infty\).

OpenStudy (solomonzelman):

so, are you good so far with why when x increases infinitely the y (the output) will also grow towards infinity?

OpenStudy (solomonzelman):

we can test this by plugging in values too; \(\large\color{black}{ \displaystyle {\rm y}=\log_{10}(x) }\) \(\large\color{black}{ \displaystyle (10,1) }\) \(\large\color{black}{ \displaystyle (100,2) }\) \(\large\color{black}{ \displaystyle (1000,3) }\) and so on.... \(\large\color{black}{ \displaystyle (10^{\rm n},{\rm n}) }\) this is growing very slowly, but it will and certain can get an infinitely large output.

OpenStudy (solomonzelman):

ask anything so far, or verify that you got everything so far. I don't want to go on without that.

OpenStudy (cutiecomittee123):

I get it so far

OpenStudy (cutiecomittee123):

@SolomonZelman

OpenStudy (solomonzelman):

Ok,

OpenStudy (solomonzelman):

Now, lets see what happens to \(\large\color{black}{ \displaystyle {\rm y}=\log(x) }\) when the value of x we are plugging in closer and closer approaches 0.

OpenStudy (solomonzelman):

What is the value of the function when x=0.1? \(\large\color{black}{ \displaystyle y= \log(x) }\) plugging 0.1 for x \(\large\color{black}{ \displaystyle y= \log(0.1) }\) re-writing as a fraction \(\large\color{black}{ \displaystyle y= \log\left(\frac{1}{10}\right) }\) now I will apply some rules that exponents have. \(\large\color{black}{ \displaystyle y= \log\left(\frac{1}{10^{1}}\right) }\) \(\large\color{black}{ \displaystyle y= \log\left(10^{-1}\right) }\) the exponent of the log, goes on the outside (you know that, right?) \(\large\color{black}{ \displaystyle y= (-1)\log\left(10\right) }\) log(10) is 1, so \(\large\color{black}{ \displaystyle y= (-1)\times 1=-1 }\) so when \(x=0.1\) then \(y=-1\)

OpenStudy (solomonzelman):

while you read this, I will post some more values, to give you an idea of the function behavior.

OpenStudy (cutiecomittee123):

Okay got it

OpenStudy (cutiecomittee123):

I knew that the exponent of the log goes on the outside before the log

OpenStudy (cutiecomittee123):

and I learned about how 10 is the number you use as a base if there is not one given with general logs and with natural logs you use e

OpenStudy (solomonzelman):

We have a point (0.1, -1). NOW, we will plug some x values that are closer to 0, and will follow the same algorithm (the same steps) to find y. what is the v-value at x=0.01? (choose 0.01 that is closer to 0, and plugged it into the function) \(\large\color{black}{ \displaystyle y= \log(0.01) }\) \(\large\color{black}{ \displaystyle y= \log\left(\frac{1}{100}\right) }\) \(\large\color{black}{ \displaystyle y= \log\left(\frac{1}{10^{2}}\right) }\) \(\large\color{black}{ \displaystyle y= \log\left(10^{-2}\right) }\) \(\large\color{black}{ \displaystyle y= (-2)\log\left(10\right) }\) \(\large\color{black}{ \displaystyle y= (-2)\times 1=-2 }\) so when \(x=0.01\) then \(y=-2\) so the points as x approaches zero so far go the following way: (0.1 , -1) (0.01 , -2)

OpenStudy (solomonzelman):

yes, glad you are getting it, and what you said is correct;)

OpenStudy (solomonzelman):

I will show more values

OpenStudy (solomonzelman):

when x=0.001, y=? \(\large\color{black}{ \displaystyle y= \log(0.001) }\) \(\large\color{black}{ \displaystyle y= \log\left(\frac{1}{1000}\right) }\) \(\large\color{black}{ \displaystyle y= \log\left(\frac{1}{10^{3}}\right) }\) \(\large\color{black}{ \displaystyle y= \log\left(10^{-3}\right) }\) \(\large\color{black}{ \displaystyle y= (-3)\log\left(10\right) }\) \(\large\color{black}{ \displaystyle y= (-3)\times 1=-3 }\) so when \(x=0.001\) then \(y=-3\) so the points as x approaches zero so far go the following way: (0.1 , -1) (0.01 , -2) (0.001 , -3)

OpenStudy (solomonzelman):

and so would be true for x=n (for any negative integer n)

OpenStudy (cutiecomittee123):

Okay

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y= \log(0.\underline{0...0}1) }\) (n zeros ^) \(\large\color{black}{ \displaystyle y= \log\left(\frac{1}{1\underline{0...0}}\right) }\) (n zeros ^) \(\large\color{black}{ \displaystyle y= \log\left(\frac{1}{10^{n}}\right) }\) \(\large\color{black}{ \displaystyle y= \log\left(10^{-n}\right) }\) \(\large\color{black}{ \displaystyle y= (-n)\log\left(10\right) }\) \(\large\color{black}{ \displaystyle y= (-n)\times 1=-3 }\) so when \(x=0.\underline{0...0}1\) then \(y=-n\) (n zeros ^) so the points as x approaches zero so far go the following way: (0.1 , -1) (0.01 , -2) (0.001 , -3) ... etc, and till \((0.\underline{00...0}1~,~-n)\) (n zeros ^)

OpenStudy (solomonzelman):

just as you can plug a super-close to 0 x-value, when x= (some tiny decimal), same way you can get a negative with a huge magnitude.

OpenStudy (cutiecomittee123):

So really its just (-infinity, +inifinity) as the range?

OpenStudy (cutiecomittee123):

what about finding the x Intercept??

OpenStudy (solomonzelman):

x intercept: set y=0, and solve for x y intercept: set x=0, and solve for y

OpenStudy (solomonzelman):

well, y-intercept wouldn't always exist .... but if it does, than this is what you got to do.